# Rethinking Graph Structure Learning in the Era of LLMs

### v3

@@ Appendix I Hyperparameter Settings
For other baselines, we primarily adopt the optimal hyperparameters reported in their original papers. If no such configurations are available, we utilize Optuna to automatically determine the best hyperparameter settings for these baselines.
@@
Reviewing the three modules of LLaTA (detailed descriptions can be found in Sec. 4.2-4.4), we observe that Step 1: Topology-aware In-context Construction (Sec. 4.2) and Step 3: Leaf-oriented Two-step Sampling (Sec. 4.4) have relatively fixed computation in which LLMs do not actively participate or play a secondary role. This is not directly related to our goal of visualizing the LLM inference through the case study to provide a deeper analysis of LLaTA’s interpretability. Therefore, we focus on Step 2: Tree-prompted LLM Inference (Sec. 4.3), specifically analyzing its three components: Reception-aware Leaf Augmentation, Community of Thought, and Leaf Dependency Allocation. In Fig. 7, we use a low-level community as an example and offer a detailed demonstration of the LLaTA inference.
@@
Appendix L Hyperparameter Analysis
@@
Figure 8: The performance of LLaTA with different $\epsilon$ on Cora and Citeseer datasets.
@@
Figure 9: The performance of LLaTA with different $\theta$ and $r$ on Child and History datasets.
To further answer Q3, we analyze the impact of the similarity threshold $\epsilon$ in Sec. 4.3. Additionally, we conducted further analysis on the impact of parameters $\theta$ and $r$ in Sec. 4.4 across more datasets.
L.1 Analysis of Hyperparameter $\epsilon$
Specifically, we evaluate the node classification performance of LLaTA under varying $\epsilon$ on the Cora and Citeseer, with the results presented in Fig. 9.
The experimental findings indicate that the hyperparameter $\epsilon$ is critical in determining the quality of the in-context information provided to LLMs and the effectiveness of their inference. A small $\epsilon$ value may lead to the aggregation of nodes from different classes, introducing noise into the in-context and thereby reducing inference accuracy. Conversely, an excessively large $\epsilon$ may result in over-filtering, disregarding important connections between similar nodes, ultimately leading to a loss of meaningful relationships and failing to provide high-quality contextual information for LLMs. Our results suggest that an appropriate range for $\epsilon$ is between 0.4 and 0.6, which balances the trade-off between minimizing noise and preserving essential connections. Specifically, the optimal $\epsilon$ value is 0.45 for Cora and 0.55 for Citeseer. In conclusion, the selection of an appropriate $\epsilon$ value plays a critical role in constructing high-quality in-context information for LLMs, ultimately enhancing the overall performance and robustness of LLaTA across diverse tasks and datasets.
L.2 Analysis of Hyperparameters $\theta$ and $r$
To thoroughly analyze the impact of $\theta$ and $r$ on LLaTA’s performance, we conducted additional experiments on the Child and History datasets, with the results presented in Fig. 9.
Our evaluation on the Child dataset shows pronounced sensitivity to the two-step sampling frequency $r$. Accuracy peaks at $r=10$ (62.28%) before declining to 61.69% at $r=25$, confirming that excessive edge additions ($r\geq 25$) disrupt heterophilic structures. The semantic sampling parameter $\theta$ maintains stable performance (61.87%-61.98%) within $[5,15]$, while extreme values ($\theta=3$ or $20$) cause 1-2% drops due to sampling limitations. The History dataset demonstrates LLaTA’s robustness to $r$ variations (84.32%-85.28% accuracy), with only 0.02% difference between $r=5$ (85.34%) and $r=25$ (85.32%). For $\theta$, peak accuracy at $\theta=10$ (85.28%) varies by less than 1% across $[5,15]$, This systematic comparison reveals that $\theta$ optimization should be performed after determining $r$, particularly for heterophilic graphs where $r$ significantly impacts structural integrity while $\theta$ mainly affects sampling quality.
These findings collectively validate the dataset-dependent nature of $r$’s impact and the universal stability of $\theta$ observed in our primary experiments. The Child dataset’s heterophily demands careful $r$ calibration to avoid overhomophilization, whereas History’s homophily permits broader parameter flexibility. Both datasets corroborate that $\theta\in[5,15]$ serves as a reliable default range, though practitioners should prioritize $r$ tuning when deploying LLaTA on heterophilic graphs. The consistent performance patterns across three distinct graph types (Cora, Child, History) strengthen the generalizability of our parameter guidelines.
Appendix M Efficiency Analysis of LLaTA
Table 8: Complete runtime(h) results of LLaTA across all 10 datasets.
@@
Method Cora Citeseer Pubmed WikiCS Instagram Reddit Ratings Child History Photo LLaTAGCN 1.593 1.974 11.872 7.081 6.850 20.136 14.583 14.016 25.394 29.373 LLaTAGAT 2.592 2.974 12.871 8.080 7.849 21.138 15.583 15.014 26.393 30.369 LLaTASAGE 3.594 3.975 13.875 9.085 8.852 22.139 16.588 16.016 27.398 31.341
@@
In this section, we present the complete runtime performance of LLaTA across 10 tagged datasets in Table 8, along with a comprehensive analysis of our method’s computational efficiency advantages.
@@
Current graph structure learning (GSL) methods face significant efficiency bottlenecks, particularly in large-scale or noisy graphs. Traditional approaches—including metric learning, probabilistic modeling, and direct optimization—often require computationally expensive pairwise similarity calculations (e.g., kernel functions or attention mechanisms) or iterative optimization of adjacency matrices[36]. These operations exhibit quadratic or higher complexity relative to node counts, limiting scalability. For LLM-based GSL Methods (e.g., GraphEdit [7], LLM4RGNN [31]), efficiency issues are further exacerbated by the inherent overhead of large-model inference and fine-tuning. Empirical results reveal that these methods exceed 72 hours on larger datasets among the 10 benchmarks (As shown in Table 1), primarily due to their reliance on full-parameter tuning and dense graph operations.
@@
Enabled by these two key innovations, our framework achieves superior efficiency by: (1) Eliminating Fine-Tuning: Bypassing LLM parameter updates reduces GPU memory and time costs by orders of magnitude. (2) Sublinear Complexity: The tree sampler operates locally within communities, avoiding global pairwise calculations. (3) Parallelizability: Community-level operations enable distributed processing, scaling to graphs with $10^{5+}$ nodes where baseline LLM-GSL methods fail.

### v4

@@ Appendix I Experimental Setup
For other baselines and backbones, we primarily adopt the optimal hyperparameters reported in their original papers. If no such configurations are available, we utilize Optuna to automatically determine the best hyperparameter settings for these baselines.
@@
Reviewing the three modules of LLaTA (detailed descriptions can be found in Sec. 4.1-4.3), we observe that Step 1: Topology-aware In-context Construction (Sec. 4.1) and Step 3: Leaf-oriented Two-step Sampling (Sec. 4.3) have relatively fixed computation in which LLMs do not actively participate or play a secondary role. This is not directly related to our goal of visualizing the LLM inference through the case study to provide a deeper analysis of LLaTA’s interpretability. Therefore, we focus on Step 2: Tree-prompted LLM Inference (Sec. 4.2), specifically analyzing its three components: Reception-aware Leaf Augmentation, Community of Thought, and Leaf Dependency Allocation. In Fig. 7, we use a low-level community as an example and offer a detailed demonstration of the LLaTA inference.
@@
Figure 8: Node classification performance on the Citeseer and WikiCS dataset under real-world scenarios (Sparsity [Edge Removal] and Noise [Edge Addition]).
Appendix L Robustness Analysis
To further examine the robustness of LLaTA in practical environments, we conduct additional experiments under sparsity and noise scenarios, where the graph structure is perturbed by randomly removing or adding edges. The results on four benchmark datasets (Cora, Pubmed, Citeseer, and WikiCS) are shown in Fig. 3 and Fig. 8.
Overall, LLaTA demonstrates strong robustness across both settings. In the sparsity scenario (edge removal), LLaTA consistently outperforms all baselines across different perturbation levels. This robustness stems from our tree-based optimization, where structural entropy guides the identification of uncertain regions, ensuring that key communities remain stable even when connections are missing. As a result, the information loss caused by sparsity is effectively mitigated.
In the noise scenario (edge addition), LLaTA also achieves competitive performance. The community-aware tree prompts allow the model to suppress the influence of noisy edges by relying more heavily on semantic consistency and hierarchical community information. Nevertheless, we observe that as the edge addition rate becomes large (e.g., 0.6 or 0.8), LLaTA’s performance gradually deteriorates. This phenomenon can be attributed to the irreversible negative effect of excessive noisy edges, which interfere with the initialization of the structural encoding tree. Importantly, although LLaTA remains among the top-performing methods, LLM4RGNN occasionally surpasses it under high noise rates. This is expected, since LLM4RGNN is explicitly designed to handle adversarial settings by leveraging fine-tuned LLMs for edge-level detection and correction.
In summary, these results highlight two key insights. First, LLaTA exhibits strong resilience to sparsity and moderate levels of noise, validating the effectiveness of its topology-aware and community-guided design. Second, the performance gap under severe noise suggests a promising direction for future work: integrating more powerful noise detection mechanisms into our tree-based optimization to further enhance robustness under adversarial or heavily corrupted graphs.
@@
Figure 9: The node classification accuracy of LLaTA with different $\epsilon$ on Cora, Citeseer, WikiCS and History datasets.
@@
Figure 10: The node classification accuracy of LLaTA with different $\theta$ and $r$ on Cora, Reddit, Child and Photo datasets.
Appendix M Hyperparameter Analysis
To further answer Q3, we provide more detailed experimental results and analysis for the four hyperparameters ($K$ in Sec. 4.1, $\epsilon$ in Sec. 4.2, $\theta$ and $r$ in Sec. 4.3).
M.1 Analysis of Hyperparameter $K$
We further examine the effect of the encoding tree height $K$, with results shown in Fig. 5 (left). The experiments reveal that the optimal $K$ varies across datasets, reflecting differences in graph structural complexity. Moderate depths ($K=2$ or $3$) generally achieve the best performance, as they strike a balance between capturing multi-level community information and avoiding redundant hierarchy. However, as $K$ increases beyond 5, accuracy consistently drops for all datasets, and the decline becomes pronounced at $K=6$. This suggests that excessively deep trees may dilute meaningful community structures and introduce noise, thereby harming the quality of structural optimization. Overall, these findings highlight that $K$ should be tuned according to graph complexity, while overly large values should be avoided due to their adverse effect on performance.
M.2 Analysis of Hyperparameter $\epsilon$
We further study the effect of $\epsilon$, which controls the threshold for constructing in-context information, on four datasets: Cora, Citeseer, WikiCS, and History, as shown in Fig. 9.
The results confirm that $\epsilon$ plays a crucial role in balancing noise reduction and information preservation. When $\epsilon$ is too small (e.g., $\epsilon=0.4$), nodes from different classes are aggregated, introducing noisy contexts and leading to reduced inference accuracy. In contrast, when $\epsilon$ becomes too large (e.g., $\epsilon=0.6$), the model over-filters, discarding meaningful connections and thereby losing valuable semantic and structural cues. Across datasets, the optimal $\epsilon$ varies slightly depending on graph characteristics: Cora and History achieve peak performance at $\epsilon=0.45$, Citeseer at $\epsilon=0.55$, and WikiCS at $\epsilon=0.5$. Despite these variations, the overall performance remains stable within the range $[0.45,0.55]$, suggesting that this interval provides a reliable trade-off between minimizing cross-class noise and preserving intra-class relationships.
In summary, $\epsilon$ is a sensitive yet well-bounded hyperparameter: choosing $\epsilon\in[0.45,0.55]$ generally yields robust performance across both homophilic and heterophilic graphs. This validates the effectiveness of our design in constructing high-quality in-context information for LLMs, while also offering practitioners practical guidance for parameter tuning in diverse real-world scenarios.
M.3 Analysis of Hyperparameters $\theta$ and $r$
To provide a comprehensive understanding of how $\theta$ (semantic sampling size) and $r$ (two-step sampling frequency) influence LLaTA, we conduct experiments on five representative datasets, with results summarized in Fig. 5 and Fig. 10.
Impact of $r$. We observe that $r$ exerts a stronger dataset-dependent effect. On heterophilic graphs such as Child, accuracy is highly sensitive to the choice of $r$: performance steadily improves up to $r=10$, but drops when $r$ is further increased (e.g., $r=25$). This degradation arises because excessive edge additions overwrite useful heterophilic patterns, forcing the structure toward artificial homophily. By contrast, on homophilic datasets such as History, Cora, and Photo, performance remains stable across a wide range of $r$ values, with fluctuations typically within 0.5%. These results indicate that while heterophilic graphs require careful calibration of $r$ to preserve structural integrity, homophilic graphs tolerate larger $r$ without substantial accuracy loss.
Impact of $\theta$. The semantic sampling parameter $\theta$ exhibits remarkable stability across all datasets. Within the range $[5,15]$, performance variations remain minor, typically under 1%. Extreme settings, however, show performance degradation: when $\theta=3$, the sampling space becomes too restrictive, limiting semantic coverage; when $\theta=20$, excessive candidate expansion introduces noise, weakening the reliability of semantic similarity. Importantly, these effects are consistent across homophilic and heterophilic datasets, demonstrating that $\theta$ primarily influences the quality of semantic sampling rather than the structural optimization itself.
From these findings, two guidelines emerge: (1) $r$ should be prioritized during tuning, especially for heterophilic graphs where improper calibration may lead to over-homogenization of node neighborhoods; (2) $\theta\in[5,15]$ can be regarded as a reliable default choice for diverse graph types, providing a balance between semantic coverage and efficiency. Taken together, these results not only validate the robustness of LLaTA across varying parameter settings but also highlight its adaptability: while $r$ controls the sensitivity to graph topology, $\theta$ primarily refines semantic discrimination. This complementary role of the two hyperparameters further underscores the stability of our framework under both homophilic and heterophilic conditions.
M.4 General Guidance on Hyperparameter Selection
Here we provide additional guidelines for selecting the four key hyperparameters in LLaTA, based on both theoretical considerations and empirical observations:
Encoding Tree Height $K$. As discussed in Sec. 4.1, the height of the encoding tree determines the number of hierarchical community divisions. Its optimal value is closely related to the graph scale and the number of node categories. Larger graphs with more categories typically exhibit more complex hierarchical structures, thus requiring deeper trees. Empirically, we observe that the optimal $K$ usually falls within the range of 2 to 5 for most datasets.
Threshold $\epsilon$. The threshold $\epsilon$ governs text propagation and aggregation. Its optimal setting depends on the quality of both textual information and the encoding tree. When either is of low quality, a higher $\epsilon$ enforces stricter filtering, thereby reducing noise and preventing unreliable information from propagating. With higher-quality inputs, smaller $\epsilon$ values can be adopted to retain richer contextual information.
Candidate Set Size $\theta$. The parameter $\theta$ controls the size of the candidate set for semantic similarity sampling. Its optimal value depends on both the quality of node text and the original graph structure. When textual attributes and structural quality are high, LLM-generated semantic embeddings are more reliable, allowing for a larger $\theta$ to improve coverage. Conversely, lower-quality data may benefit from smaller $\theta$ values to reduce noise.
Sampling Frequency $r$. The frequency $r$ in two-step sampling is influenced by the average degree of the original and optimized graphs. For graphs with higher average degrees, larger $r$ values are generally needed to preserve sufficient structural connectivity. In contrast, for sparse graphs, moderate $r$ values are preferable to avoid excessive edge additions.
These guidelines offer practical insights for tuning LLaTA across diverse graph types, providing a balance between general applicability and dataset-specific adaptability.
Table 8: Complete runtime(h) results of LLaTA across all $1$ datasets.
@@
Method Cora Citeseer Pubmed WikiCS Instagram Reddit Ratings Child History Photo ArXiv LLaTAGCN 1.593 1.974 12.873 7.081 7.850 21.136 15.583 15.016 26.394 30.372 53.175 LLaTAGAT 1.598 1.980 12.881 7.124 7.861 21.145 15.589 15.025 26.451 30.468 53.479 LLaTASAGE 1.583 1.958 12.861 7.076 7.847 21.131 15.578 15.009 26.489 30.369 53.177
Appendix N Efficiency Analysis of LLaTA
@@
In this section, we present the complete runtime performance of LLaTA across 11 tagged datasets in Table 8, along with a comprehensive analysis of our method’s computational efficiency advantages.
@@
Current graph structure learning (GSL) methods face significant efficiency bottlenecks, particularly in large-scale or noisy graphs. Traditional approaches—including metric learning, probabilistic modeling, and direct optimization—often require computationally expensive pairwise similarity calculations (e.g., kernel functions or attention mechanisms) or iterative optimization of adjacency matrices[40]. These operations exhibit quadratic or higher complexity relative to node counts, limiting scalability. For LLM-based GSL Methods (e.g., GraphEdit [7], LLM4RGNN [34]), efficiency issues are further exacerbated by the inherent overhead of large-model inference and fine-tuning. Empirical results reveal that these methods exceed 72 hours on larger datasets among the 11 benchmarks (As shown in Table 1), primarily due to their reliance on full-parameter tuning and dense graph operations.
@@
Furthermore, we provide an efficiency-oriented acceleration scheme for large-scale graphs to further enhance the scalability of our method. To reduce computational overhead while maintaining the effectiveness of structure optimization, we introduce a composite scoring mechanism to identify the most needed low-level communities for optimization.
$\text{Score}(\mathcal{C}^{\ell})=\lambda_{1}\cdot\mathcal{H}_{\text{struct}}(\mathcal{C}^{\ell})+\lambda_{2}\cdot\mathcal{H}_{\text{label}}(\mathcal{C}^{\ell})+\lambda_{3}\cdot\Big(1-\mu_{\text{textsim}}(\mathcal{C}^{\ell})\Big),$ (41)
where $\mathcal{H}_{\text{struct}}(\mathcal{C}^{\ell})$ denotes the structural entropy, quantifying topological complexity; $\mathcal{H}_{\text{label}}(\mathcal{C}^{\ell})$ measures label distribution entropy, capturing heterogeneity; and $\mu_{\text{textsim}}(\mathcal{C}^{\ell})$ is the mean pairwise semantic similarity within the community. Thus, $1-\mu_{\text{textsim}}(\mathcal{C}^{\ell})$ reflects semantic inconsistency, assigning higher scores to communities with lower internal similarity. By ranking communities with this score, we prioritize optimization and sampling on top-ranked candidates. This targeted strategy significantly reduces computational overhead while preserving the effectiveness of structure optimization, enabling LLaTA to scale to large datasets such as ArXiv within practical time limits. In our experiments on the arXiv dataset, we applied the community sampling algorithm to select 40% of the communities for Tree-prompted LLM Inference process, which significantly improved efficiency on large-scale graphs while preserving the effectiveness of our method.
(c) Efficiency-Oriented Community Selection: The composite scoring mechanism adaptively identifies and optimizes only the most informative communities, reducing redundant computation and ensuring scalability to large graphs.
Enabled by these design choices, our framework achieves superior efficiency through the following three aspects: (1) Eliminating Fine-Tuning: By completely bypassing parameter updates of LLMs, our method reduces both GPU memory footprint and training time by orders of magnitude. This makes the framework applicable to resource-constrained environments where fine-tuning large models is infeasible. (2) Sublinear Complexity: The tree-based sampler operates locally within communities rather than across the entire graph, effectively reducing global pairwise computations. This results in sublinear complexity with respect to the number of nodes, making our method scalable to large and dense graphs. (3) Parallelizability and Efficiency-Oriented Community Selection: Community-level operations are inherently parallelizable and can be distributed across multiple processors, enabling scalability to graphs with $10^{5+}$ nodes where prior LLM-GSL methods fail. Moreover, our composite scoring mechanism further enhances scalability by adaptively identifying and optimizing only the most promising communities, thereby reducing redundant computation while maintaining the effectiveness of structure optimization (e.g., on the arXiv dataset we optimize only 40% of the communities while achieving competitive performance).